# ALMA FACTS III. High-Resolution CO(2-1)/CO(1-0) Maps of Twelve Nearby Galaxies

### v1

@@ 5 Discussion
This work presents high-resolution maps of a sample of nearby spiral galaxies, an early result from the ALMA-FACTS survey. While we plan on more quantitative analyses in our subsequent papers, the maps already offer the first opportunity to compare the variations of amongst galaxies. The maps appear to show some general trends with respect to galactic morphology. While discussions on morphology tend to be empirical, we attempt to characterize the general trends in terms of galactic structures. In particular, the galactic centers, bars, and spiral arms appear to play their own distinct roles in the organization of the molecular gas conditions traced by , in addition to their mass distributions as often discussed in the literature. These empirical characterizations also seem consistent with the results in recent case studies (e.g., den Brok et al., 2021; Egusa et al., 2022; Maeda et al., 2022; den Brok et al., 2023).
@@
We will first explain an idea for the stellar and gas morphology-based classifications and supporting dynamical models in the literature (section 5.1). We then discuss the behaviors of in the galactic centers, bars, bar ends, and spiral arms, as well as radial and azimuthal variations of along this empirical classification (section 5.2).
5.1 Morphology-Based Classification
@@
For characterizing general trends in , we rearrange the sample galaxies based on (1) their optical morphologies and (2) molecular gas morphologies (i.e., molecular gas distributions) with emphasis on bar and spiral arm structures. Their optical morphologies are classified as barred (SB and SAB) and unbarred spiral galaxies (SA), which appear to primarily determine the behaviors of across the disks. For barred galaxies, there is a suggested picture on the gas evolutionary sequence based on hydrodynamical simulations and observations (Sakamoto et al., 1999; Sheth et al., 2005; Yu et al., 2022), which appears to help understand the behaviors of amongst the barred galaxies. We summarize this evolutionary sequence and discuss our arrangement of the sample galaxies.
5.1.1 Bar-Driven Gas Transport
@@
In the bar-driven gas transport model, the stellar bar potential organizes the gas motions to form the offset ridges on the leading side of the bar. The interactions between the gas along the offset ridges (e.g., shocks) and/or with the stellar potential (e.g., torque) cause the gas to lose angular momentum, and the gas falls towards the center from the offset ridges (Wada, 1994; Sheth et al., 2002). This picture is supported by observations by Sakamoto et al. (1999) and Sheth et al. (2005). They studied the central regions of barred spiral galaxies, and found that the central kiloparsec of barred spiral galaxies have higher concentrations of molecular gas than their unbarred counterparts. They attributed the concentrations of molecular gas in their central regions to the bar-driven gas transport by stellar bars to the centers.
5.1.2 Classification
@@
Figure 11 shows the maps from Section 4 arranged primarily by their optical morphological classifications (SB, SAB, or SA), and secondarily by their gas morphology for the barred galaxies. The secondary part is based on the presence or absence of gas along the offset ridges. Even when the offset ridges are not present, all of the barred galaxies have central condensations of gas, and thus fit within the bar-driven gas transport picture, if the absence of the ridges is due to the gas inflow. Hence, using the optical and gas morphological classifications, we arrange the maps (ordered by galaxy number) starting from barred galaxies with shorter offset ridges, i.e., lacking molecular gas between the central region and spiral arms, to barred galaxies with longer offset ridges and having clearer spiral arms extending from the bar ends, and then to the unbarred galaxies.
@@
In the arranged evolutionary sequence of Figure 11, the barred galaxies NGC 1512, NGC 3351, NGC4579, and NGC 4536 (top row) have shorter offset ridges and/or show an absence of gas between the central regions and their surrounding spiral arm structures. Following these, the barred galaxies NGC 1097, NGC 3627, and NGC 4321 (left middle row) show longer offset ridges and have clearer spiral arms extending out of the bar ends. The rest of the galaxies are unbarred (the bottom row and the rightmost in the middle row). Figure 13 is the same as Figure 11, but shows the galaxies on the same physical scale. NGC 4826 (unbarred) is placed succeeding the barred galaxies (middle row) since it has a central gas condensation/region with a linear size more comparable to those of the barred galaxies (e.g., Section 4.1 and Figure 13).
In Table 3, we include the derived average molecular gas surface densities from within () and outside () of the fiducial 1 kpc central region. We assume the Milky-Way = 4.35 M⊙pc-2 value (Bolatto et al., 2013). The are consistent with Sakamoto et al. (1999) and Sheth et al. (2005) for common galaxies after accounting for differences in the adopted CO-to-H2 conversion factor and observational uncertainties. Amongst the barred galaxies, those with longer offset ridges have overall higher molecular gas surface densities in the central 1 kpc region and the rest of the disk than those with shorter offset ridges.
@@ 5.2 Structural Dependence of
We discuss the overall trends in seen amongst the barred and unbarred galaxies, and then within each galactic structure (center, bar, bar ends, and spiral arms). Figure 12 shows the radial distributions (each pixel) of , in the same arrangement as Figure 11. The binned means are shown in red and the secondary x-axis (top) is in arcseconds.
@@
Overall, the barred galaxies (top and left middle rows) follow a general trend: is high in the center ( 1.0), low along the bar (, increases at the bar ends ( 0.7), and then gradually declines to lower towards the disk outskirts, with fluctuations in the spiral arms. The radial distributions of in the unbarred galaxies (right middle and bottom rows) look similar to the outer parts of the barred galaxies (Komugi et al., 2025).
@@
There are also some variations in the azimuthal direction in the sample galaxies. However, due to the sensitivity limitations, the interarm regions of the galaxies are not readily detected (Figure 11). Variations in the azimuthal direction can be seen as scatters at each radius in the radial distributions. With these trends in mind, we discuss as a function of structure in the following sections.
5.2.1 Central Regions: High
@@
It has been known that barred galaxies often have high gas concentrations in their central regions (Sakamoto et al., 1999; Sheth et al., 2005; Yu et al., 2022). The same central gas concentrations are apparent in the barred galaxies in our observations (Figures 3-4). In addition, the sample of the unbarred spiral galaxies do not show such gas concentrations at the centers besides NGC 4826. This study (section 4.1) shows that the central 1-2 kiloparsec regions within barred galaxies also comprise of high/very-high ratio gas, and that the gas physical conditions change within the central gas condensations. The elevated line-ratios in their galactic centers are clear in the radial distributions.
@@
The presence of high-ratio gas across these central regions indicates that the molecular gas there has higher and and is thus denser and/or warmer than the typical molecular gas conditions, (, )=(300 cm-3, 10 K), in the disks.
@@
The high ratios may be the result of dynamical processes and/or on-going star formation in the galaxy centers. As a dynamical process, the high ratios in the central regions of the barred galaxies across the sample could be due to the bar-driven gas transport to the centers, discussed in section 5.1.1. The influence of the bar may compress the gas (increase in gas density) and/or trigger star formation in galactic centers (increase in gas kinetic temperature by stellar feedback) and lead to the high-ratios. The enhancement of star formation in the centers of barred galaxies has been studied e.g., in Ho et al. (1997).
@@ 5.2.2 Along the Bar: Low
The lower suggests that and remain low in the bar, even when the gas is concentrated along the offset ridges. This may explain why star formation efficiency is often low in the bar regions of galaxies (e.g., Downes et al., 1996; Koda & Sofue, 2006; Momose et al., 2010; Maeda et al., 2023). Recently, Rodríguez et al. (2024) analyzed the 3.3 m polycyclic aromatic hydrocarbon (PAH) emission in 19 nearby star-forming galaxies to identify star clusters in their early dust-embedded phases. They also found that, in general, there is an absence of PAH emitters in bars (including our FACTS sample: NGC 1512, NGC 3351, NGC 3627, and NGC 4321).
The generally low along the offset ridges may also be important in considering gas dynamics within the bar. It has been suggested that the offset ridges are regions of strong shocks (Athanassoula, 1992a, b; Sellwood & Wilkinson, 1993), which should naturally increase and , and hence, . Instead, the gas concentration along the offset ridges may be caused by the mere crowding of gas orbits with only occasional collisions (Wada, 1994; Koda & Sofue, 2006), which could keep low in general. In the case of orbit crowding, the gas falls towards the center by the loss of angular momentum due to the gravitational torque of the stellar bar potential, as the gas resides longer on the leading side of the bar. This gas infall process takes longer than the rapid gas transport due to the strong shocks. In addition to explaining the low along the bar, this cloud-orbit model may also explain the frequency of offset ridges in the sample of barred galaxies; along the evolutionary sequence, the presence and frequency of the offset ridges (from long to short to none) may imply the molecular offset ridges should last long (Section 5.1.1).
5.2.3 Bar Ends: High
@@
The bar ends also show higher spatially and in the radial distributions () (Figures 11-12). The high ratios again suggest that the molecular gas located there has relatively higher and/or .
@@
5.2.4 Spiral Arms
Spiral arms are also where the gas builds up. fluctuates between low and high values along the molecular gas spiral arms (Figures 11-12). It is remarkable that the fluctuations appear similar between the spiral arms in the barred galaxies and those of the unbarred galaxies. Most of the high-ratio gas in these regions appear to be associated with HII regions, and thus, star formation. This has been found in Galactic and extragalactic studies (e.g., Sakamoto et al., 1994; Seta et al., 1998; Koda et al., 2012, 2020; Egusa et al., 2022; Koda et al., 2025).
@@ 5.2.4 Spiral Arms
In the analysis of the nearby barred spiral galaxy M83, which had higher resolution and sensitivity, Koda et al. (2025) recently found that changes systematically from a low ratio in the interarm regions to high ratio in the bar and spiral arms even without star formation. They concluded that the large-scale galactic dynamics plays an important role in the evolution of molecular gas, in addition to the local impacts of stellar feedback. Indeed, within the spiral arms of NGC 4321, appears to change from low to high, when comparing one side of the arm to the other. This may also apply to NGC 1097 and NGC 3627.
5.2.5 by Structure
@@
clearly varies within each galaxy. The dependence on galactic structures (center, bar, bar ends, and spiral arms) is apparent in Section 4.1. For a more quantitative comparison, we define regions to enclose the structures in each galaxy (Figure 11). We created masks for the center, bar, and bar ends. Our adopted region definitions are not meant to be exact boundaries, but are an attempt to capture the rough areas of the structures of interest. We describe the definitions and distribution in each region below.
@@
For the central regions, we adopt either 1 or 2 kpc diameter areas around the galactic center (projected in the sky). We adopt the fiducial 1 kpc for all unbarred galaxies. For the barred galaxies, we showed that the size of the high-ratio region in the center sometimes extends beyond this in section 4.1. We adopt 2 kpc for all barred galaxies except for NGC3351 and NGC3627.
@@
For the bar and bar ends, we define regions based on the cataloged bar structure (cyan ellipse) in Herrera-Endoqui et al. (2015), which is based on Spitzer 3.6 m from S4G (Sheth et al., 2010; S4G Team, 2020). We use their on-sky PAbar, semi-major axis , and ellipticity (see Table 1). The bar region is enclosed by the cyan ellipse but does not include the bar ends (magenta). The bar ends are 2 kpc circles centered on the “tips” of the ellipse. We adopt the definitions of the spiral arms from Querejeta et al. (2021). The area outside the center, bar, bar end, and spiral arms is analyzed as interarm regions.
Figure 14 shows the distributions based on these definitions. They reflect the general trends discussed in section 5.2. Central regions of barred galaxies show high . Bars show mostly low . Bar ends for the galaxies with relatively gas-rich offset ridges typically show high . The other galaxies with gas-poor offset ridges lack the gas to be measured, or tend to exhibit lower ratios than those of the gas-rich barred galaxies. Interarm regions contain mostly low-ratio gas and show overall lower than the spiral arms.
@@
6 Systematic Trends and Comparison to Literature
In sections 4 and 5, we showed that systematically varies as a function of galactic structure in the twelve nearby spiral galaxies. Such structural variations were not previously clear from single-dish surveys (den Brok et al., 2021; Yajima et al., 2021; Leroy et al., 2022). Komugi et al. (2025, Paper II) conducted the analysis using the ALMA-TP component only. They were not able to resolve the galactic structures with their coarser resolution, but found that most of the barred galaxies showed steeper radial gradients, whereas all the unbarred galaxies showed flat or shallower gradients.

### v2

@@ V Discussion
This work presents high-resolution maps of a sample of nearby spiral galaxies, an early result from the ALMA-FACTS survey. While we plan on more quantitative analyses in our subsequent papers, the maps already offer the first opportunity to compare the variations of amongst galaxies at high-resolution ( 100-200 pc). The maps appear to show some general trends with respect to galactic morphology. While, in general, discussions on morphology tend to be empirical, we attempt to characterize the general trends in terms of galactic structures. In particular, the galactic centers, bars, and spiral arms appear to play their own distinct roles in the organization of the molecular gas conditions traced by , in addition to their mass distributions as often discussed in the literature. These empirical characterizations also seem consistent with the results in recent case studies at high-resolution ( 100-200 pc) (e.g., den Brok et al., 2021; Egusa et al., 2022; Maeda et al., 2022; den Brok et al., 2023)
@@
We will first explain an idea for the stellar and gas morphology-based classifications and supporting dynamical models in the literature (section V.1). We then discuss the behaviors of in the galactic centers, bars, bar ends, and spiral arms, as well as radial and azimuthal variations of along this empirical classification (section V.2).
V.1 Morphology-Based Classification
@@
For characterizing general trends in , we rearrange the sample galaxies based on (1) their optical morphologies and (2) molecular gas morphologies (i.e., molecular gas distributions) with emphasis on bar and spiral arm structures. Their optical morphologies are classified as barred (SB and SAB) and unbarred spiral galaxies (SA), which appear to primarily determine the behaviors of across the disks. For barred galaxies, there is a suggested picture on the gas evolutionary sequence based on hydrodynamical simulations and observations (Sakamoto et al., 1999; Sheth et al., 2005; Yu et al., 2022), which appears to help understand the behaviors of amongst the barred galaxies. We summarize this sequence and discuss our arrangement of the sample galaxies.
V.1.1 Bar-Driven Gas Transport
@@
In the bar-driven gas transport model, the stellar bar potential organizes the gas motions to form the offset ridges on the leading side of the bar. The interactions between the gas along the offset ridges (e.g., shocks) and/or with the stellar potential (e.g., torque) cause the gas to lose angular momentum, and the gas falls towards the center from the offset ridges (Wada, 1994; Sheth et al., 2002). This picture is supported by observations. Sakamoto et al. (1999) and Sheth et al. (2005) studied the central regions of barred spiral galaxies, and found that the central kiloparsec of barred spiral galaxies have higher concentrations of molecular gas than their unbarred counterparts. They, and others in the literature, attributed the concentrations of molecular gas in their central regions to the bar-driven gas transport by stellar bars to the centers (Sakamoto et al., 1999; Sheth et al., 2005; García-Burillo et al., 2009; Sormani et al., 2023). In this process, the amount of gas in the bar should decrease, if no gas is inflowing from outside the bar.
V.1.2 Classification
@@
Figure 11 shows the maps from Section IV arranged primarily by their optical morphological classifications (SB, SAB, or SA), and secondarily by their gas morphology for the barred galaxies. The secondary part is based on the presence or absence of gas along the offset ridges. Even when the offset ridges are not present, all of the barred galaxies have central condensations of gas, and thus fit within the bar-driven gas transport picture, if the absence of the ridges is due to the gas inflow. Hence, using the optical and gas morphological classifications, we arrange the maps (ordered by galaxy number) starting from barred galaxies with shorter offset ridges, i.e., lacking molecular gas between the central region and spiral arms, to barred galaxies with longer offset ridges and having clearer spiral arms extending from the bar ends, and then to the unbarred galaxies.
@@
In the arranged gas morphological sequence of Figure 11, the barred galaxies NGC 1512, NGC 3351, NGC4579, and NGC 4536 (top row) have shorter offset ridges and/or show an absence of gas between the central regions and their surrounding spiral arm structures. Following these, the barred galaxies NGC 1097, NGC 3627, and NGC 4321 (left middle row) show longer offset ridges and have clearer spiral arms extending out of the bar ends. The rest of the galaxies are unbarred (the bottom row and the rightmost in the middle row). Figure 13 is the same as Figure 11, but shows the galaxies on the same physical scale. NGC 4826 (unbarred) is placed succeeding the barred galaxies (middle row) since it has a central gas condensation/region with a linear size more comparable to those of the barred galaxies (e.g., Section IV.1 and Figure 13).
The barred galaxies have higher average surface brightness in their central regions than the unbarred galaxies. If we simply assume the Milky-Way CO-to-H2 conversion factor of = 4.35 M⊙pc-2 (Bolatto et al., 2013), then this translates to higher average molecular gas surface densities in the centers of the barred galaxies (Table 3). This is consistent with Sakamoto et al. (1999) and Sheth et al. (2005), which was interpreted as an outcome of bar-driven gas transport towards the center. We note the conversion of surface brightness to molecular mass has the caveat that may vary especially in the galaxy centers (e.g., Sandstrom et al., 2013; Teng et al., 2023).
@@ V.2 Structural Dependence of
We discuss the overall trends in seen amongst the barred and unbarred galaxies, and then within each galactic structure (center, bar, bar ends, and spiral arms). Figure 12 shows the radial distributions in two ways: as scatter plots to show the sampling (i.e., data density; top) and as density plots with the same data normalized by the peak number of points per radial bin to show the radial trend (bottom). They are arranged in the same order as Figure 11. The binned means (area-weighted) are shown in blue and orange for the barred galaxies with short or long ridges respectively, and teal for unbarred galaxies. The secondary x-axis is in arcseconds, and the blue dotted line indicates the bar length.
@@
Overall, the barred galaxies (top and left middle rows) follow a general trend: is high in the center ( 1.0), low along the bar (, increases at the bar ends ( 0.7), and then lowers beyond the bar ends or flattens, with fluctuations in the spiral arms. The higher ratios in the bar ends are remarkable, given that the bar end areas are only in a small azimuthal range but still appear as high in these azimuthally-averaged plots. The radial distributions of in the unbarred galaxies (right middle and bottom rows) look similar to the outer parts of the barred galaxies (Komugi et al., 2025).
@@
There are also some variations in the azimuthal direction in the sample galaxies. However, due to the sensitivity limitations, most interarm regions of the galaxies are not readily detected (Figure 11). Variations in the azimuthal direction can be seen as scatters at each radius in the radial distributions. With these trends in mind, we discuss as a function of structure in the following sections.
V.2.1 Central Regions: High
@@
It has been known that barred galaxies often have high gas concentrations in their central regions (Sakamoto et al., 1999; Sheth et al., 2005; Yu et al., 2022). The same central gas concentrations are apparent in the barred galaxies in our observations (Figures 3-4). In addition, the sample of the unbarred spiral galaxies do not show such gas concentrations at the centers besides NGC 4826. This study (section IV.1) shows that the central 1-2 kiloparsec regions within barred galaxies also comprise of high/very-high ratio gas, and that the gas physical conditions change within the central gas condensations. is elevated significantly and consistently in the centers, which has also been found in earlier studies, e.g. Braine & Combes (1992); Leroy et al. (2009); den Brok et al. (2021). The elevated line-ratios in their galactic centers are clear in the radial distributions.
@@
The presence of high-ratio gas across these central regions indicates that the molecular gas there has higher and and is thus denser and/or warmer than the typical molecular gas conditions, (, )=(300 cm-3, 10 K), in the disks (Koda et al., 2012).
@@
The high ratios may be the result of dynamical processes and/or on-going star formation in the galaxy centers. As a dynamical process, the high ratios in the central regions of the barred galaxies across the sample could be due to the bar-driven gas transport to the centers, discussed in section V.1.1. The influence of the bar may compress the gas (increase in gas density) and/or trigger star formation in galactic centers (increase in gas kinetic temperature by stellar feedback) and lead to the high-ratios. The enhancement of star formation in the centers of barred galaxies has been studied e.g., in Ho et al. (1997).
@@ V.2.2 Along the Bar: Low
The lower suggests that and remain low in the bar (Koda et al., 2012), even when the gas is concentrated along the offset ridges. This may explain why star formation efficiency is often low in the bar regions of galaxies (e.g., Downes et al., 1996; Koda & Sofue, 2006; Momose et al., 2010; Maeda et al., 2023). Recently, Rodríguez et al. (2024) analyzed the 3.3 m polycyclic aromatic hydrocarbon (PAH) emission in 19 nearby star-forming galaxies to identify star clusters in their early dust-embedded phases. They also found that, in general, there is an absence of PAH emitters in bars (including our FACTS sample: NGC 1512, NGC 3351, NGC 3627, and NGC 4321).
The generally low along the offset ridges may also be important in considering gas dynamics within the bar. It has been suggested that the offset ridges are regions of strong shocks (Athanassoula, 1992a, b; Sellwood & Wilkinson, 1993), which should naturally increase and , and hence, . Instead, the gas concentration along the offset ridges may be caused by the mere crowding of gas orbits with only occasional collisions (Wada, 1994; Koda & Sofue, 2006), which could keep low in general. In the case of orbit crowding, the gas falls towards the center by the loss of angular momentum due to the gravitational torque of the stellar bar potential, as the gas resides longer on the leading side of the bar (Wada, 1994). This gas infall process would take longer than the rapid gas infall by the strong shocks, which could make offset ridges live longer, and appear more frequently in a sample of barred galaxies. In the future, it may be interesting to check the frequency of the offset ridges (from long to short to none), in a larger sample.
V.2.3 Bar Ends: High
@@
The bar ends also show higher spatially and in the radial distributions often ) (Figures 11-12). This is most apparent in the galaxies with longer offset ridges (see also section V.2.5). The high ratios again suggest that the molecular gas located there has relatively higher and/or (Koda et al., 2012). When there is not much gas in the bar region (as in the short-offset ridge galaxies) it is difficult to assess. NGC 1512 shows an increase towards the bar ends in its radial profile, but this is unclear in NGC 3351 and NGC 4579. NGC 4536 is affected by projection effects (section IV.1.10).
@@
V.2.4 Spiral Arms
Spiral arms are also where the gas builds up. fluctuates between low and high values along the molecular gas spiral arms (Figures 11-12) (see also e.g., Druard et al., 2014; Koda et al., 2025). It is remarkable that the fluctuations appear similar between the spiral arms in the barred galaxies and those of the unbarred galaxies. High-ratio gas in these regions sometimes appear to be associated with HII regions, and thus, star formation. This has been found in Galactic and extragalactic studies (e.g., Sakamoto et al., 1994; Seta et al., 1998; Koda et al., 2012, 2020; Egusa et al., 2022; Koda et al., 2025).
@@ V.2.4 Spiral Arms
In the analysis of the nearby barred spiral galaxy M83, which had higher resolution and sensitivity, Koda et al. (2025) recently found that changes systematically from a low ratio in the interarm regions to high ratio in the bar and spiral arms even without star formation. They concluded that the large-scale galactic dynamics plays an important role in the evolution of molecular gas, in addition to the local impacts of stellar feedback. In this sample, NGC 1097, NGC 3627, and NGC 4321 show more coherent molecular spiral arms, with some parts possibly showing a similar transition from low to high .
V.2.5 by Structure
@@
clearly varies within each galaxy. The dependence on galactic structures (center, bar, bar ends, and spiral arms) is apparent in Section IV.1. For a more quantitative comparison, we define regions to enclose the structures in each galaxy (Figure 11). We created masks for the center, bar, and bar ends. Our adopted region definitions are not meant to be exact boundaries, but are an attempt to capture the rough areas of the structures of interest. We describe the definitions and distribution in each region below.
@@
For the central regions, we adopt either 1 or 2 kpc diameter areas around the galactic center (projected in the sky). We adopt the fiducial 1 kpc for all unbarred galaxies. For the barred galaxies, we showed that the size of the high-ratio region in the center sometimes extends beyond this in section IV.1. We adopt 2 kpc for all barred galaxies except for NGC3351 and NGC3627.
@@
For the bar and bar ends, we define regions based on the cataloged bar structure (cyan ellipse) in Herrera-Endoqui et al. (2015), which is based on Spitzer 3.6 m from S4G (Sheth et al., 2010; S4G Team, 2020). We use their on-sky PAbar, semi-major axis , and ellipticity (see Table 1). The bar region is enclosed by the cyan ellipse but does not include the bar ends (magenta). The bar ends are 2 kpc circles centered on the “tips” of the ellipse.
We adopt the definitions of the spiral arms from Querejeta et al. (2021). The area outside the center, bar, bar end, and spiral arms is analyzed as interarm regions.
Using these definitions, we compared the distributions for adjacent structure pairs (center-bar, bar-bar end, and bar end-spiral arm) and the spiral arm-interarm regions by performing the Kolmogorov-Smirnov (KS) two-sample test using the ks_2samp function in scipy. The null hypothesis is that the two samples originate from the same parent distribution. We applied this test including only every fifth pixel (Nyquist sampling) (1) within each galaxy and (2) all galaxies combined. We find p-values (majority lower than ) meaning the null hypothesis can be rejected in all the structure pairs compared in all galaxies, except in two cases: bar-bar end in NGC 1512 (p-value 0.4) and bar end-spiral arm in NGC 4536 (p-value 0.3). In NGC 1512 there is not much gas in the bar, and in NGC 4536 the bar-end definition is ambiguous due to its projection (section IV.1.10). (Note: with full-beam sampling, the NGC1097 bar-bar end and NGC4321 bar end-spiral arm comparisons also show p-values 0.05. However, p-values are still in most cases, and the overall discussion does not change.)
Figure 14 shows the corresponding distributions. They reflect the general trends discussed in section V.2. Central regions of barred galaxies show high . Bars show mostly low . Bar ends for the galaxies with relatively longer offset ridges more often show high . We note that the longer offset ridged galaxies contribute the most in panels (c-d) because more emission is detected there (see section V.2.3).
The KS test also suggests the arm-interarm distributions are different (p-values ). The histogram for the interarms is slightly but systematically shifted towards lower (Figure 14). The contrast could be higher if the sensitivity is higher because low brightness/density gas, expected more in the interarms, tends to have lower . We note that our equal flux-based cut in CO(1-0) and CO(2-1) (SN5) practically corresponds to a higher density threshold for CO(2-1) if is low (Table 2, column (6)).
@@
VI Systematic Trends and Comparison to Literature
In sections IV and V, we showed that systematically varies as a function of galactic structure in the twelve nearby spiral galaxies. Such structural variations were not previously clear from single-dish surveys (den Brok et al., 2021; Yajima et al., 2021; Leroy et al., 2022). Komugi et al. (2025, Paper II) conducted the analysis using the ALMA-TP component only. They were not able to resolve the galactic structures with their coarser resolution, but found that most of the barred galaxies showed steeper radial gradients, whereas all the unbarred galaxies showed flat or shallower gradients.